Prove that √(1+√3) is irrational, assuming that √3 is irrational

1
Expert's answer
2022-02-02T10:51:35-0500

Solution:

Given, "\sqrt3" is irrational.

To prove that "\sqrt{1+\sqrt3}" is irrational.

Proof:

Assume "\sqrt{1+\sqrt3}" is rational, so that "\sqrt{1+\sqrt3}=\dfrac ab", where a,b are co-primes and "b\ne0" .

"\Rightarrow 1+\sqrt3=\dfrac {a^2}{b^2}
\\ \Rightarrow \sqrt3=\dfrac {a^2}{b^2}+1"

On left side, "\sqrt3" is irrational as it is given.

On right side, "\dfrac {a^2}{b^2}+1" must be rational because "\dfrac ab" is rational by assumption.

Now, an irrational cannot be equal to rational.

We get a contradiction.

So, our assumption is wrong.

Thus,"\sqrt{1+\sqrt3}" is irrational.